Ziggy's Famous Yogurt blends regular yogurt that is 3% fat with its no-fat yogurt to obtain low fat yogurt that is 1% fat. How many pounds of regular yogurt and how many pounds of no-fat yogurt should be mixed to obtain 60 pounds of low-fat yogurt?: Ziggy's Famous Yogurt blends regular yogurt that is 3% fat with its no-fat yogurt to obtain low fat yogurt that is 1% fat. How many pounds of regular yogurt and how many pounds of no-fat yogurt should be mixed to obtain 60 pounds of low-fat yogurt?
Answer by josmiceli(2035)

Call the pounds of regular yogurt you need r
Call the pounds of no-fat yogurt you need n
You have to end up with a mixture that is 1% fat.
Stated in words this is:
(pounds of fat you end up with) / (total pounds of mixture) = % fat
What is the pounds of fat you end up with?
There's no fat from the no-fat yogurt
The pounds of fat from the regular yogurt is .03r
The problem says you need to end up with 60 pounds of yogurt
.03r / 60 = .01
.03r = .6
r = 20
That leaves 60 - 20 = 40 pounds of the no-fat that need
to be added
20 pounds of the regular and 40 pounds of the no-fat are needed
check answer:
.03*20 / 60 = .01
.6/60 = .01
.6 = .6
OK